# An Efficient Local Search Approach for Polarized Community Discovery in Signed Networks

### v1

@@ Abstract
Signed networks, where edges are labeled as positive or negative to indicate friendly or antagonistic interactions, offer a natural framework for studying polarization, trust, and conflict in social systems. Detecting meaningful group structures in these networks is crucial for understanding online discourse, political division, and trust dynamics. A key challenge is to identify groups that are cohesive internally yet antagonistic externally, while allowing for neutral or unaligned vertices. In this paper, we address this problem by identifying $k$ polarized communities that are large, dense, and balanced in size. We develop an approach based on Frank-Wolfe optimization, leading to a local search procedure with provable convergence guarantees. Our method is both scalable and efficient, outperforming state-of-the-art baselines in solution quality while remaining competitive in terms of computational efficiency.
@@ 1 Introduction
Signed networks extend traditional graph representations by associating each edge with a positive or negative number, indicating friendly or antagonistic relationships. Originating from studies on social dynamics in the 1950s (Harary, 1953), signed networks introduce fundamental differences in graph structure that make many algorithms designed for unsigned networks inapplicable (Tang et al., 2016; Bonchi et al., 2019; Tzeng et al., 2020). These challenges have fueled extensive research in recent years, leading to advances in signed network embeddings, signed clustering, and signed link prediction. We refer to the survey by (Tang et al., 2016) for a comprehensive review of these methods. Most relevant to this paper is the problem of signed clustering, which we split into two categories: (i) signed network partitioning (SNP), and (ii) polarized community discovery (PCD). The latter is the problem studied in this paper.
@@
The goal of signed clustering is to identify $k$ conflicting groups (clusters) where intra-cluster similarity is maximized (predominantly positive) and inter-cluster similarity is minimized (predominantly negative). This problem has numerous real-world applications (Tang et al., 2016), particularly in social networks, where vertices represent individuals and edges capture friendly or antagonistic relationships (e.g., shared or opposing political views). Detecting conflicting groups in such networks is crucial for analyzing polarization (Adamic and Glance, 2005; Yardi and Boyd, 2010; Xiao et al., 2020), echo chambers (Garrett, 2009; Flaxman et al., 2016), and the spread of misinformation (Shu et al., 2017; Cooke, 2018; Yang et al., 2019).
@@
In the SNP problem, the $k$ groups must form a partition of the vertices, meaning every vertex must be included. Spectral methods based on the signed Laplacian have been widely used to tackle this problem (Kunegis et al., 2010; Chiang et al., 2012; Mercado et al., 2019; Cucuringu et al., 2019). Alternatively, formulating SNP explicitly as an optimization problem leads to the well-studied correlation clustering (CC) problem (Bansal et al., 2004), which is known to be APX-hard. Consequently, numerous approximation algorithms have been developed (Bansal et al., 2004; Charikar et al., 2005; Demaine et al., 2006; Ailon et al., 2008), with local search methods standing out for their strong performance in both clustering quality and computational efficiency (Thiel et al., 2019; Chehreghani, 2023; Aronsson and Chehreghani, 2024a; b).
@@
The problem formulation of PCD is identical to that of SNP, except that the $k$ clusters are not required to form a partition of the vertices, allowing some vertices to remain unassigned. The goal is therefore to only find the dense subgraphs of polarized communities. This accounts for cases where certain vertices are neutral w.r.t. the underlying conflicting group structure. For example, in a social network with a heated political debate, many users may not engage in the dispute, and their interactions might not align with any specific faction. There is a substantial body of work addressing this problem, but most approaches focus on identifying only two communities (Bonchi et al., 2019; Xiao et al., 2020; Ordozgoiti et al., 2020; Fazzone et al., 2022; Niu and Sariyüce, 2023; Gullo et al., 2024). As a result, they do not easily generalize to arbitrary $k$. To our knowledge, only two works specifically tackle PCD for arbitrary $k$. (Chu et al., 2016) formulated the task as a constrained quadratic optimization problem and proposes an efficient algorithm that iteratively refines small subgraphs, avoiding the costly computation of the full adjacency matrix. (Tzeng et al., 2020) introduced a spectral method based on maximizing a discrete Rayleigh quotient, which extends the seminal work of (Bonchi et al., 2019) to accommodate arbitrary $k$. Their method is known to produce highly imbalanced communities in terms of size (Gullo et al., 2024).
The main contributions of this paper are as follows: (i) We propose a novel formulation for the PCD problem that encourages more balanced communities, addressing a key limitation of (Tzeng et al., 2020). (ii) We demonstrate that this formulation is particularly well-suited for optimization via the block-coordinate Frank-Wolfe algorithm (Frank and Wolfe, 1956; Lacoste-Julien et al., 2013). (iii) We establish a simple yet effective equivalent local search method with a provable convergence rate. (iv) We propose techniques that allow the local search method to scale to large networks. (v) Finally, through extensive experiments, we show that our approach consistently outperforms state-of-the-art baselines.
@@ 2 Problem Formulation
We start by introducing the relevant notation, followed by an introduction to correlation clustering (CC), which is connected to our problem. Finally, we present the problem of polarized community discovery (PCD).
2.1 Notation
Consider a signed network $G=(V,E)$, where $V$ is the set of objects and $E$ the set of edges. The weight of an edge $(i,j)\in E$ is represented by the element $A_{i,j}\in\{-1,0,+1\}$ of an adjacency matrix $A$. The matrix $A$ is symmetric with zeros on the diagonal, which means $A_{i,j}=A_{j,i}$ and $A_{i,i}=0$. We use $A_{i,:}$ and $A_{:,j}$ to denote row $i$ and column $j$ of $A$, respectively. While we restrict all similarities to be in $\{-1,0,+1\}$ (for clarity), all methods presented in the paper extend to arbitrary similarities in $\mathbb{R}$. We can decompose the adjacency matrix as $A=A^{+}-A^{-}$ where $A^{+}=\max(A,0)$ and $A^{-}=\max(-A,0)$. A clustering with $k$ clusters is denoted $S_{[k]}=\{S_{1},\dots,S_{k}\}$, where each $S_{m}\subseteq V$ is the set of objects assigned to cluster $m\in[k]=\{1,\dots,k\}$. Let $N^{+}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{+}_{i,j}$ and $N^{-}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{-}_{i,j}$ be the sum of positive and absolute negative intra-cluster similarities, respectively. Furthermore, let $N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{i\in S_{m}}\sum_{j\in S_{p}}A^{+}_{i,j}$ and $N^{-}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{i\in S_{m}}\sum_{j\in S_{p}}A^{-}_{i,j}$ be the sum of positive and absolute negative inter-cluster similarities, respectively.
@@ 2.2 Correlation Clustering
We begin by noting that for CC, unlike PCD to be discussed in the next subsection, a clustering $S_{[k]}$ is a partition of $V$, meaning $V=\bigcup_{m\in[k]}S_{m}$ and each $S_{m}$ is disjoint. A notable feature of CC is its ability to automatically determine the number of clusters based on the similarities encoded in $A$. However, a variant of this problem has also been studied where the number of clusters, $k$, is specified as an input (Giotis and Guruswami, 2006). Automatic detection of the number of clusters could be a desirable property of a clustering algorithm. However, constraining the number of clusters to $k$ can act as a form of regularization, and has been shown to produce higher-quality clusters in many scenarios (Chehreghani et al., 2012). Given this, the $k$-CC problem can be defined as shown below.
@@
In other words, we want to find a clustering that maximizes the sum of intra-cluster similarities and minimizes the sum of inter-cluster similarities. The following proposition presents alternative objectives equivalent to maximizing Eq. 1. While this is known in the CC literature (Chehreghani, 2013), we include a complete summary here to better contextualize our problem.

### v2

@@ Abstract
Signed networks—where edges are labeled as positive or negative to represent friendly or antagonistic interactions—offer a natural framework for analyzing polarization, trust, and conflict in social systems. Detecting meaningful group structures in such networks is crucial for understanding online discourse, political divisions, and trust dynamics. A key challenge is to identify communities that are internally cohesive and externally antagonistic, while allowing for neutral or unaligned vertices. In this paper, we propose a method for identifying $k$ polarized communities that addresses a major limitation of prior methods: their tendency to produce highly size-imbalanced solutions. We introduce a novel optimization objective that avoids such imbalance. In addition, it is well known that approximation algorithms based on local search are highly effective for clustering signed networks when neutral vertices are not allowed. We build on this idea and design the first local search algorithm that extends to the setting with neutral vertices while scaling to large networks. By connecting our approach to block-coordinate Frank-Wolfe optimization, we prove a linear convergence rate, enabled by the structure of our objective. Experiments on real-world and synthetic datasets demonstrate that our method consistently outperforms state-of-the-art baselines in solution quality, while remaining competitive in computational efficiency.
@@ 1 Introduction
Signed networks extend traditional graph representations by associating each edge with a positive or negative number, indicating friendly or antagonistic relationships. Originating from studies on social dynamics in the 1950s (10.1307/mmj/1028989917, ), signed networks introduce fundamental differences in graph structure that make many algorithms designed for unsigned networks inapplicable (DBLP:journals/csur/TangCAL16, ; DBLP:conf/cikm/BonchiGGOR19, ; DBLP:conf/nips/TzengOG20, ). These challenges have fueled extensive research in recent years, leading to advances in signed network embeddings, signed clustering, and signed link prediction. We refer to the survey by (DBLP:journals/csur/TangCAL16, ) for a comprehensive review of these methods. Most relevant to this paper is the problem of signed clustering, which we split into two categories: (i) signed network partitioning (SNP), and (ii) polarized community discovery (PCD). The latter is the problem studied in this paper.
@@
The goal of signed clustering is to identify $k$ clusters where intra-cluster similarity is maximized (predominantly positive) and inter-cluster similarity is minimized (predominantly negative). This problem has numerous real-world applications (DBLP:journals/csur/TangCAL16, ), particularly in social networks, where vertices represent individuals and edges capture friendly or antagonistic relationships (e.g., shared or opposing political views). Detecting conflicting groups in such networks is crucial for analyzing polarization (polarization, ; polarizationtwitter, ; DBLP:conf/www/XiaoOG20, ), echo chambers (echo, ; 10.1093/poq/nfw006, ), and the spread of misinformation (DBLP:journals/sigkdd/ShuSWTL17, ; alma997565503602341, ; DBLP:conf/aaai/YangSWG0019, ).
@@
In the SNP problem, the $k$ groups must form a partition of the vertices, meaning every vertex must be included. Spectral methods based on the signed Laplacian have been widely used to tackle this problem (DBLP:conf/sdm/KunegisSLLLA10, ; DBLP:conf/cikm/ChiangWD12, ; DBLP:conf/icml/MercadoT019, ; DBLP:conf/aistats/CucuringuDGT19, ). Alternatively, formulating SNP explicitly as an optimization problem leads to the well-studied correlation clustering (CC) problem (DBLP:journals/ml/BansalBC04, ), which is known to be APX-hard. Consequently, numerous approximation algorithms have been developed (DBLP:journals/ml/BansalBC04, ; CharikarGW05, ; DemaineEFI06, ; AilonCN08, ), with local search methods standing out for their strong performance in both clustering quality and computational efficiency (DBLP:conf/aaai/ThielCD19, ; Chehreghani22_shift, ; DBLP:journals/tmlr/AronssonC24, ; aronsson2024informationtheoreticactivecorrelationclustering, ).
@@
The problem formulation of PCD is identical to that of SNP, except that the $k$ clusters are not required to form a partition of the vertices, allowing some vertices to remain unassigned. The goal is therefore to only find the dense subgraphs of polarized communities. This accounts for cases where certain vertices are neutral w.r.t. the underlying conflicting group structure. For example, in a social network with a heated political debate, many users may not engage in the dispute, and their interactions might not align with any specific faction. There is a substantial body of work addressing this problem, but most approaches focus on identifying only two communities (DBLP:conf/cikm/BonchiGGOR19, ; DBLP:conf/www/XiaoOG20, ; DBLP:conf/www/OrdozgoitiMG20, ; DBLP:journals/pvldb/FazzoneLDTB22, ; DBLP:conf/www/NiuS23, ; DBLP:journals/ml/GulloMT24, ). As a result, they do not easily generalize to arbitrary $k$. To our knowledge, only two works specifically tackle PCD for arbitrary $k$. (DBLP:conf/kdd/ChuWPWZC16, ) formulated the task as a constrained quadratic optimization problem and proposes an efficient algorithm that iteratively refines small subgraphs, avoiding the costly computation of the full adjacency matrix. (DBLP:conf/nips/TzengOG20, ) introduced a spectral method based on maximizing a discrete Rayleigh quotient, which extends the seminal work of (DBLP:conf/cikm/BonchiGGOR19, ) to accommodate arbitrary $k$. These methods are known to produce highly imbalanced communities in terms of size (DBLP:journals/ml/GulloMT24, ).
The main contributions of this paper are as follows: (i) We propose a novel formulation for the PCD problem that encourages more balanced communities, addressing a key limitation of previous work that typically optimize polarity (DBLP:conf/nips/TzengOG20, ) (see Eq. 2). As demonstrated in our experiments, optimizing polarity often leads to clustering solutions with multiple empty clusters. The importance of promoting balanced communities (to avoid trivial solutions where all objects are placed in one or few clusters) is well established in the graph clustering literature DBLP:conf/cikm/ChiangWD12 . In Appendix E, we expand on this, and provide some examples of practical scenarios where (reasonably) balanced communities are favorable in our context. We note that DBLP:journals/ml/GulloMT24 also proposes an objective for PCD called $\gamma$-polarity aimed at addressing cluster imbalance; however, it is restricted to the case of $k=2$ clusters. In contrast, our objective supports an arbitrary number of clusters $k$. Nonetheless, we compare to their method experimentally for $k=2$, and explain how it differs conceptually in Appendix F. (ii) Motivated by the effectiveness of local search-based approximation algorithms for CC (and many other machine learning models), we propose the first scalable local search algorithm for PCD, which explicitly allows for neutral objects. (iii) We establish a linear convergence rate of our local search algorithm by connecting it to block-coordinate Frank-Wolfe optimization (fw, ; DBLP:conf/icml/Lacoste-JulienJSP13, ). This connection is made possible utilizing the specific structure of our proposed optimization objective. (iv) We propose techniques that allow the local search method to scale to large networks. (v) Finally, through extensive experiments on commonly used real-world and synthetic datasets in previous work on PCD, we show that our approach consistently outperforms state-of-the-art baselines in terms of (a) recovering ground-truth solutions and (b) finding high quality solutions of reasonable cluster size balance.
@@ 2 Problem Formulation
We start by introducing the relevant notation, followed by an introduction to CC, which is connected to our problem. Finally, we describe PCD, including our novel formulation of the problem.
Notation. Consider a signed network $G=(V,E)$, where $V$ is the set of objects and $E$ the set of edges. The weight of an edge $(i,j)\in E$ is represented by the element $A_{i,j}\in\{-1,0,+1\}$ of an adjacency matrix $A$. The matrix $A$ is symmetric with zeros on the diagonal, which means $A_{i,j}=A_{j,i}$ and $A_{i,i}=0$. We use $A_{i,:}$ and $A_{:,j}$ to denote row $i$ and column $j$ of $A$, respectively. While we restrict all similarities to be in $\{-1,0,+1\}$ (for clarity), all methods presented in the paper extend to arbitrary similarities in $\mathbb{R}$. We can decompose the adjacency matrix as $A=A^{+}-A^{-}$ where $A^{+}=\max(A,0)$ and $A^{-}=\max(-A,0)$. A clustering with $k$ clusters is denoted $S_{[k]}=\{S_{1},\dots,S_{k}\}$, where each $S_{m}\subseteq V$ is the set of objects assigned to cluster $m\in[k]=\{1,\dots,k\}$. Let $N^{+}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{+}_{i,j}$ and $N^{-}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{-}_{i,j}$ be the sum of positive and absolute negative intra-cluster similarities, respectively. Furthermore, let $N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{i\in S_{m}}\sum_{j\in S_{p}}A^{+}_{i,j}$ and $N^{-}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{i\in S_{m}}\sum_{j\in S_{p}}A^{-}_{i,j}$ be the sum of positive and absolute negative inter-cluster similarities, respectively.
@@ 2.1 Correlation Clustering
We begin by noting that for CC, unlike PCD to be discussed in the next subsection, a clustering $S_{[k]}$ is a partition of $V$, meaning $V=\bigcup_{m\in[k]}S_{m}$ and each $S_{m}$ is disjoint. A notable feature of CC is its ability to automatically determine the number of clusters DBLP:conf/kdd/BonchiGL14 , but here we we focus on the $k$-constrained variant of CC (DBLP:journals/toc/GiotisG06, ) as it is most relevant to our problem. The $k$-CC problem can be defined as shown below.
@@
In other words, the goal is to find a clustering that (i) maximizes intra-cluster similarities and (ii) minimizes inter-cluster similarities. In the CC literature, it is known that maximizing certain subsets of terms in Eq. 1—such as the total number of agreements, $N^{+}_{\text{intra}}+N^{-}_{\text{inter}}$—is equivalent to maximizing the full objective ethz-a-010077098 . However, this equivalence does not hold for PCD when neutral objects are allowed, as each of the four terms in Eq. 1 contributes uniquely to the decision of whether an object should be clustered or left neutral. We formally show this in Appendix A, and thus focus on the full objective in the next section when we introduce PCD. CC is known to be NP-hard DBLP:journals/ml/BansalBC04 , and many approximation algorithms have been proposed (DBLP:journals/ml/BansalBC04, ; CharikarGW05, ; DemaineEFI06, ; AilonCN08, ), with local search methods standing out for their strong performance in both clustering quality and computational efficiency (DBLP:conf/aaai/ThielCD19, ; Chehreghani22_shift, ; DBLP:journals/tmlr/AronssonC24, ).
2.2 Polarized Community Discovery
For PCD, we introduce a neutral set $S_{0}$. As a result, each object in $V$ is either assigned to one of the non-neutral clusters $S_{1},\dots,S_{k}$ or designated as neutral by placing it in $S_{0}$. Consequently, a clustering $S_{[k]}$ is no longer a partition of $V$ and we have $S_{0}=V\setminus\bigcup_{m\in[k]}S_{m}$ (although all clusters are still disjoint). Given this, the goal of PCD is to identify non-neutral clusters $S_{[k]}$ such that (i) a large value of the objective in Eq. 1 is obtained (consistent with CC), and (ii) the graph induced by the non-neutral objects is as dense as possible (i.e., most edge weights are $+1$ or $-1$). Any object that hinders either of these goals should be assigned to the neutral set $S_{0}$. This includes ambiguous objects, such as those with significant similarity to multiple clusters or those with inherently weak associations (e.g., low-degree nodes). Importantly, there exists a natural trade-off between the size and density of the non-neutral clusters: small clusters can trivially achieve high density. As we will demonstrate, our objective allows for a flexible balance of this trade-off.
In prior work, it is common to encourage the presence of neutral objects by penalizing large/sparse non-neutral clusters. This is typically done by normalizing Eq. 1 by the number of non-neutral objects, i.e.,
$\frac{(N^{+}_{\text{intra}}-N^{-}_{\text{intra}})+\alpha(N^{-}_{\text{inter}}-N^{+}_{\text{inter}})}{\sum_{m\in[k]}|S_{m}|},$ (2)
where $\alpha\in\mathbb{R}$ is used to balance (i) maximization of intra-cluster similarities and (ii) minimization of inter-cluster similarities. If $\alpha=1/(k-1)$, Eq. 2 is commonly referred to as polarity in prior work and is a well-established objective for PCD (DBLP:conf/cikm/BonchiGGOR19, ; DBLP:conf/nips/TzengOG20, ). This choice of $\alpha$ was proposed in DBLP:conf/nips/TzengOG20 , based on the observation that the number of intra-similarities scale linearly with $k$, while the number of inter-similarities grow quadratically. This choice prevents inter-similarities from dominating the objective. We use this value of $\alpha$ throughout the paper unless otherwise stated.
However, as highlighted in (DBLP:journals/ml/GulloMT24, ) (and in our experiments), maximizing polarity often results in highly imbalanced clustering solutions (often with multiple empty clusters). In particular, clustering solutions with the same polarity can differ significantly in terms of cluster size balance. A concrete example illustrating this issue is provided in Appendix D. (DBLP:journals/ml/GulloMT24, ) proposes a new objective called $\gamma$-polarity that addresses this issue for the special case of $k=2$. Our proposed objective is different from $\gamma$-polarity and is applicable with any arbitrary $k$. In Appendix F, we compare $\gamma$-polarity with our proposed objective, defined in Eq. 3. We also compare to (DBLP:journals/ml/GulloMT24, ) in our experiments.
In this paper, we propose an alternative objective that, instead of normalizing by the number of non-neutral objects, incorporates a regularization term by subtracting the sum of squared sizes of the non-neutral clusters.
The third term in Eq. 3 has been previously applied to the minimum cut objective for unsigned networks (Chehreghani22_shift, ). In our context (i.e., for the PCD problem), the objective in Eq. 3 achieves two goals simultaneously: it penalizes the formation of (i) large/sparse and (ii) highly imbalanced non-neutral clusters. The second property is easy to see, as for a clustering with $k$ clusters and $n$ objects, the term $\sum_{m\in[k]}|S_{m}|^{2}$ is minimized when each cluster is assigned $\frac{n}{k}$ objects (i.e., the clusters are perfectly balanced). We introduce regularization as an additive term rather than a normalization for two key reasons: (i) It allows a flexible trade-off between the number of non-neutral objects and the density of the graph induced by them (controlled by the parameter $\beta\in\mathbb{R}$), which is a desirable property in this context as discussed in the beginning of this section. This possibility is absent in the existing methods that are based on Eq. 2 (polarity). (ii) It enables the development of an efficient optimization procedure based on local search with strong convergence guarantees. In the context of CC, local search–based approximation algorithms are known to significantly outperform other methods (DBLP:conf/aaai/ThielCD19, ; Chehreghani22_shift, ). In this paper, we develop the first scalable local search algorithm specifically tailored to the PCD setting. In Section 4, we demonstrate across a range of real-world and synthetic datasets that the advantages of local search optimization (e.g., for CC) carry over to PCD as well. We are now ready to formally state our problem, and we subsequently highlight its computational complexity in Thm. 1.
Problem 2 ($k$-PCD).
Theorem 1.
Problem 2 (i.e., $k$-PCD) is NP-hard.
All proofs can be found in Appendix B.
3 Algorithms
Thm. 1 underscores the necessity of approximate methods to solve Problem 2. In this section, we demonstrate how it can be solved using Frank-Wolfe (FW) optimization fw . Specifically, we consider a variant called block-coordinate FW, which we begin by describing in the next subsection. After this, we establish its equivalence to a straightforward and provably efficient local search procedure. Next, we analyze the convergence rate of this approach. Following that, we propose practical enhancements to improve scalability, enabling the method to handle large problems. A detailed discussion of the impact of $\alpha$ and $\beta$ is deferred to Appendix C.
3.1 Block-Coordinate Frank-Wolfe Optimization
The Frank–Wolfe (FW) algorithm is one of the earliest methods for nonlinear constrained optimization fw . In recent years, it has regained popularity, particularly in machine learning, due to its scalability DBLP:conf/icml/Jaggi13 . In this paper, we use a variant of this method called block-coordinate FW DBLP:conf/icml/Lacoste-JulienJSP13 . This method yields a significantly faster optimization procedure while enjoying similar theoretical guarantees. Block-coordinate FW is applied to problems where the feasible domain can be split into blocks $\mathcal{D}=\mathcal{D}^{(1)}\times\dots\times\mathcal{D}^{(n)}\subseteq\mathbb{R}^{d}$, where each $\mathcal{D}^{(i)}\subseteq\mathbb{R}^{d_{i}}$ is convex and compact and we have $d=\sum_{i=1}^{n}d_{i}$. Let ${\bm{x}}_{[n]}$ denote the concatenation of the variables ${\bm{x}}_{i}\in\mathcal{D}^{(i)}$ from all blocks $i\in[n]$. The optimization problem is then
where $f$ is a differentiable function with an $L$-Lipschitz continuous gradient. This approach is particularly effective when optimizing $f$ w.r.t. the variables in a single block (while keeping other blocks fixed) is simple and efficient. This turns out to be the case for our problem, as will be discussed in the remainder of this section. The method is outlined in Alg. 1, where $\nabla_{i}f({\bm{x}}_{[n]})$ represents the gradient of $f({\bm{x}}_{[n]})$ with respect to block ${\bm{x}}_{i}$. When the problem involves only a single block ($n=1$), Alg. 1 reduces to the standard FW algorithm. We now show how our problem can be turned into an instance of Eq. 4. Below, we show an alternative way of writing our objective in Eq. 3.